OpenStudy (anonymous):

I need help with solving equations. I'm not sure what to do after a certain step. For example: First I did: 2 ( 3x + 5 ) = x + 8 + 5x Than: 6x + 10 = x + 8 + 5x. After that I don't know what to do. Do I subtract the 8? Or do I subtract the 5x? Can someone explain in steps please?!

OpenStudy (anonymous):

you should add and subtract numbers with numbers and variables with each other ..it's better that move numbers at one side of equation and move variables to the other side and then subtract them and you will solve it like this 6x+10=x+8+5x move number to the right side of equation: 6x-x-5x=8-10 0=-2 so equation don't have any result..

OpenStudy (anonymous):

Thank you! I appreciate your help.

OpenStudy (anonymous):

your welcome